# T-GVC: Trajectory-Guided Generative Video Coding at Ultra-Low Bitrates

### v3

@@ 3.4 Video reconstruction
Inspired by (Xiao et al. 2024; Zhang et al. 2025), we utilize motion trajectories to guide the generation process in a training-free manner by imposing trajectory-aware constraints during the denoising diffusion process. In (Xiao et al. 2024; Zhang et al. 2025), the trajectories are extracted from the feature maps of U-Net blocks in VDM. However, these approaches incur substantial computational overhead during inference, particularly for generative models with a large number of parameters. Moreover, since the motion of generated results is globally controlled, it is difficult to guide the model’s output with fine-grained precision. In contrast, T-GVC directly imposes constraint on the latent noise, leveraging the insight that the latent space in video diffusion contains crucial structural information and semantic correlation. As shown in Figure 3, the region with the highest similarity in Figure 3(a) corresponds to the position of the red point in Figure 3(c), which belongs to the same trajectory as the red point in Figure 3(a) or Figure 3(b).
For each clip of the target video with length $L$, the pipeline on the decoder side begins with decoding compressed keyframes and associated motion trajectories. After that, the keyframes $\{{\hat{K}}_{i},{\hat{K}}_{i+1}\}$ are projected into latent space via VAE (Kingma, Welling et al. 2013) encoder and concatenated with the initial noisy latent $z_{T}=\{z_{T}^{0},\ z_{T}^{1},\ \ldots,\ z_{T}^{L-1}\}$ as conditions, providing the key textual and structural features of the original video.
@@
We could update the output $\epsilon_{\theta}$ of the model in each time step:
@@
where $f(\cdot)$ is the map from noise latent $z_{t}$ to the clean latent $z_{0}$.
@@
Most existing VDMs are trained to generate videos of fixed length (e.g., 16 frames). In contrast, our framework accommodates variable intervals between adjacent keyframes. To this end, we further design a variable-length generation scheme to enhance the flexibility of pre-trained VDMs.
Specifically, for each video clip with length $L\leq 16$, we first interpolate the sparse trajectories $\tau$ to $\hat{\tau}$ with length $\hat{L}=16$ and assign corresponding positional markers to the original trajectories. The denoised latent frames at these marked positions are then selected and concatenated to form the final video. For each video clip with length $L>16$, we perform a dual-stage generation strategy where the trajectories are divided into two segments, with each segment guiding one generation stage separately. Figure 4 demonstrates the operational scenario of our framework in processing a video clip beyond 16 frames, with a theoretical maximum capacity of 30 frames. Notably, increasing the number of processing stages facilitates extended video generation with enhanced temporal coherence, while linearly escalating computational demands.
@@ 4.1 Experimental setup
Following (Wan, Zheng, and Fan 2025), we select HEVC Class B, C (Bossen 2010) as well as UVG (Mercat, Viitanen, and Vanne 2020) and MCL-JCV (Wang et al. 2016) as the test dataset to evaluate the rate-distortion (R-D) performance. All test videos are resized to $512\times 320$ and consist of 96 frames with the original frame rate. For the calculation of rate-distortion metric, the compression rate is quantified in bits per pixel (bpp), while the distortion is measured using LPIPS (Zhang et al. 2018) for perceptual quality and CLIP-SIM (Radford et al. 2021) for semantic similarity. Lower LPIPS and higher CLIP-SIM at the same bitrate represent better coding performance.
@@
We select the pre-trained DynamiCrafter (Xing et al. 2024a) video diffusion model as our inter-frame decoder. Alternative video generation models designed for open-domain image interpolation or morphing are also compatible with our framework. We remove text prompts from the original model and solely utilize trajectories and keyframes to guide the generation of inter-frames. The number of DDIM denoising steps for each video and the classifier guide scale $s\left(t\right)$ are set to 10 and $30\cdot\sqrt{1-\alpha_{t}}$, respectively. For keyframe compression, we adopt MS-ILLM (Muckley et al. 2023) under ’quality 1’ and ’quality 2’ settings. In addition, given that the video resolution is downsampled to $64\times 40$ by a factor of 8 after projected into latent space, the grid size for sparse motion sampling is configured as 64 to preserve structural consistency. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ 4.2 Quantitative and qualitative results
We compare T-GVC with traditional codecs (H.265 (Sullivan et al. 2012) and H.266 (Bross et al. 2021)) and one of the most competitive open-source neural video codec (DCVC-FM (Li, Li, and Lu 2024)) with a wide bitrate range to evaluate its coding performance at ULB. The comparison results with other generative video coding schemes are provided and discussed in Section 4.3. For H.265, we use HM-18.0 with QP = 51, 45 and 39. For H.266, we use VTM-23.7 with QP = 63, 57, 51, 45 and 39. For DCVC-FM, we set the $q$ indexes to 0, 8, 16, 24 and 32, which control the compression level. All codecs above are tested under Low Delay P configuration with intra-period==-1.
The R-D curves for the test video classes are presented in Figure 5, where the bitrate of our T-GVC is controlled by adjusting the quality of keyframes. As observed, T-GVC outperforms both traditional codecs and the neural video codec in terms of perceptual reconstruction quality and semantic fidelity at ULB conditions. Moreover, T-GVC achieves lower bitrate points (below 0.005 bpp) than GAN- or VQ-VAE-based methods (Yang, Timofte, and Van Gool 2022; Qi et al. 2025), with detailed comparisons provided in the appendix. Figure 6 shows the visual results for the sequence ‘BasketballDrill_832x480_50’ from the HEVC Class C dataset. To ensure a fair comparison, reconstructed sequences from different methods are selected at comparable bitrates. It is observed that the reconstructed videos generated by T-GVC demonstrate superior fidelity in the background regions and key foreground objects (e.g., basketball and the players in red). In contrast, VTM exhibits noticeable blocking artifacts while DCVC-FM suffers from significant high-frequency detail loss, leading to degraded texture quality. It is worth noting that although some texture and structural details in the reconstructed frames might be partially missing or inaccurate to some extent (e.g., the direction of the ball and the appearance of players in green), the overall visual quality remains satisfactory for the ULB scenarios.
@@ 4.3 Ablation study
To further verify the rationality of our design, we conduct ablation studies on two core components of the T-GVC framework: the guidance mechanism and the motion sampler module, as well as the keypoint quantity within each trajectory instance. Video sequences characterized by large motions, including $ParkScene$, BasketballDrillBasketballDrill, PartyScenePartyScene, $videoSRC02$ and $videoSRC05$ from HEVC Class B, Class C and MCL-JCV, are selected for evaluation. The coding efficiency is assessed using the Bjøntegaard Delta rate (BD-rate) (Bjontegaard 2001), where PSNR is replaced by 1LPIPS1-LPIPS. Negative BD-rate values indicate bitrate saving for equivalent perceptual quality. We apply DynamiCrafter with empty prompts (i.e., only guided by keyframes) as the comparison anchor in our experiments. The reported BD-rate results are calculated by averaging all test sequences.
@@
Sparse $K_{max}$ = 5 W/o motion bitrate -5.94
W/ motion bitrate 4.12
$K_{max}$ = 10 W/o motion bitrate -13.84
W/ motion bitrate -1.86
$\mathbf{\boldsymbol{K_{max}}}$ = 15 W/o motion bitrate -16.84
@@
$K_{max}$ = 20 W/o motion bitrate -10.58
W/ motion bitrate 5.82
Table 4: Ablation study on the keypoint quantity.
Ablation study on the guidance mechanism
In prior generative video coding methods (Zhang et al. 2024b; Wan, Zheng, and Fan 2025), text is commonly adopted as motion guidance. To demonstrate the superiority of our guidance mechanism, we implement three variants of generative frameworks based on DynamiCrafter: one guided by text (denoted as Text), one guided by trajectories (denoted as Trajectory) and the other by both text and trajectories (denoted as Text+Trajectory). Trajectories are sampled through our proposed sparse motion sampler, while text descriptions are generated from the source video using Hunyuan-Large (Sun et al. 2024) without limitation on length. The comparison results are presented in Table 2, where ”W/o motion bitrate” refer to excluding the bitrate consumed by motion-related information when calculating BPP, while ”W/ motion bitrate” indicate that this bitrate is included. It is observed that the trajectory-guided approach consistently achieves better coding performance under both evaluation conditions. In particular, the perceptual quality improvement brought by trajectory guidance makes up for the additional bitrate overhead, offering a distinct advantage over text-based guidance. When overhead is not considered, the trajectory-guided method still outperforms the text-guided method. Combining text and trajectory guidance may enhance texture on simple-motion sequences. Some visual examples in Figure 7 further demonstrate that trajectory guidance can perform more precise motion control, especially in dynamic areas. Text guidance can improve perceptual quality to some extent and combining text and trajectory guidance may enhance texture on simple-motion sequences. However, in more complex cases, text prompts may introduce artifacts that conflict with trajectory-aligned motion, leading to degraded results. For instance, as shown in the first column of Figure 7, the basketball appears behind the net, which violates the physical plausibility.
Ablation study on the motion sampler module
To validate the effectiveness of the semantic-aware motion sampling pipeline proposed in T-GVC, we further conduct an ablation study on the motion sampler module. The ablation results are presented in Table 3. Specifically, ”Sparse” refers to motion trajectories extracted with our semantic-aware motion sampling pipeline. ”Dense” denotes applying dense trajectories as guidance without sampling. ”Random” corresponds to randomly selecting sparse trajectories with $k$ keypoints ($K<K_{max}$) for each motion instance. Although dense trajectories can provide more precise motion representation for video reconstruction, such excessive bitrate overhead is intolerable for the video coding task. In contrast, our motion sampling pipeline achieves significant bitrate reduction in inter-frame motion modeling, while preserving perceptual quality comparable to that of dense trajectory guidance. Compared with a random sampling strategy, our approach thoroughly considers the impact of each motion instance on the video semantics, effectively reducing the bitrate while retaining key motion information.
Ablation study on the keypoint quantity
5 Conclution
In this paper, we propose T-GVC, a novel Trajectory-guided Generative Video Coding framework designed for ULB scenarios. Unlike existing methods that rely heavily on domain-specific priors or high-level text guidance, T-GVC introduces a semantic-aware sparse motion sampling strategy, extracting motion trajectories based on semantic importance to retain key temporal information with minimal bitrate overhead. To further enhance motion realism, a training-free latent guidance mechanism based on trajectory-aligned loss is integrated into the diffusion process, enabling accurate and physically plausible reconstructions. Experimental results demonstrate that T-GVC significantly outperforms traditional codecs and state-of-the-art generative compression methods in terms of both reconstruction quality and motion control precision.

### v4

@@ 3.3 Video reconstruction
Inspired by (xiao2024video; zhang2025training), we utilize motion trajectories to guide the generation process in a training-free manner by imposing trajectory-aware constraints during the denoising diffusion process. In (xiao2024video; zhang2025training), the trajectories are extracted from the feature maps of U-Net blocks in VDM. However, these approaches incur substantial computational overhead during inference, particularly for generative models with a large number of parameters. Moreover, since the motion of generated results is globally controlled, it is difficult to guide the model’s output with fine-grained precision. In contrast, T-GVC directly imposes constraint on the latent noise, leveraging the insight that the latent space in video diffusion contains crucial structural information and semantic correlation. As shown in Figure 3, the region with the highest similarity in Figure 3(a) corresponds to the position of the red point in Figure 3(c), which belongs to the same trajectory as the red point in Figure 3(a) or Figure 3(b).
For each clip of the target video with length $L$, the pipeline on the decoder side begins with decoding compressed keyframes and associated motion trajectories. After that, the keyframes $\{{\hat{K}}_{i},{\hat{K}}_{i+1}\}$ are projected into latent space via VAE (kingma2013auto) encoder and concatenated with the initial noisy latent $z_{T}=\{z_{T}^{0},\ z_{T}^{1},\ \ldots,\ z_{T}^{L-1}\}$ as conditions, providing the key textual and structural features of the original video.
@@
We can further update the output $\epsilon_{\theta}$ of the model in each time step via the gradient of $\mathcal{L}_{m}$:
@@
where $f(\cdot)$ is the map from noise latent $z_{t}$ to the clean latent $z_{0}$, and a detailed theoretical explanation of our proposed guidance mechanism is detailed in the appendix.
@@
Most existing VDMs are trained to generate videos of fixed length (e.g., 16 frames). To this end, we further design a variable-length generation scheme to enhance the flexibility of pre-trained VDMs. Specifically, for each video clip with length $L\leq 16$, we first interpolate the sparse trajectories $\tau$ to $\hat{\tau}$ with length $\hat{L}=16$ and assign corresponding positional markers to the original trajectories. The denoised latent frames at these marked positions are then selected and concatenated to form the final video. For each video clip with length $L>16$, we perform a dual-stage generation strategy where the trajectories are divided into two segments, with each segment guiding one generation stage separately.
@@ 4.1 Experimental setup
Following (wan2025m3), we select HEVC Class B, C (bossen2010common) as well as UVG (mercat2020uvg) and MCL-JCV (wang2016mcl) as the test dataset to evaluate the rate-distortion (R-D) performance. All test videos are resized to $512\times 320$ and consist of 96 frames with the original frame rate. For the calculation of R-D metric, the compression rate is quantified in bits per pixel (bpp), while the distortion is measured using LPIPS (zhang2018unreasonable) for perceptual quality and CLIP-SIM (radford2021learning) for semantic similarity. Lower LPIPS and higher CLIP-SIM at the same bitrate represent better coding performance.
@@
We select the pre-trained DynamiCrafter (xing2024dynamicrafter) video diffusion model as our inter-frame decoder due to its superior quality for open-domain image animation. We remove text prompts from the original model and solely utilize trajectories and keyframes to guide the generation of inter-frames. The number of DDIM denoising steps for each video and the classifier guide scale $s\left(t\right)$ are set to 10 and $30\cdot\sqrt{1-\alpha_{t}}$, respectively. For keyframe compression, we adopt MS-ILLM (muckley2023improving) under ‘quality 1’ and ‘quality 2’ settings. In addition, given that the video resolution is downsampled to $64\times 40$ by a factor of 8 after projected into latent space, the grid size for sparse motion sampling is configured as 64 to preserve structural consistency. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ 4.2 Quantitative and qualitative results
We compare T-GVC with traditional codecs (H.265 (sullivan2012overview) and H.266 (bross2021overview)) and one of the most competitive open-source neural video codec (DCVC-FM (li2024neural)) with a wide bitrate range to evaluate its coding performance at ULB. The comparison results with other generative video coding schemes are provided and discussed in Section 4.3. For H.265, we use HM-18.0 with QP = 51, 45 and 39. For H.266, we use VTM-23.7 with QP = 63, 57, 51, 45 and 39. For DCVC-FM, we set the $q$ indexes to 0, 8, 16, 24 and 32, which control the compression level. All codecs above are tested under Low Delay P configuration with intra-period==-1.
The R-D curves for the test video classes are presented in Figure 4, where the bitrate of our T-GVC is controlled by adjusting the quality of keyframes. As observed, T-GVC outperforms both traditional codecs and the neural video codec in terms of perceptual reconstruction quality and semantic fidelity at ULB conditions. Moreover, benefiting from powerful generative priors, T-GVC achieves lower bitrate points (below 0.005 bpp) than GAN- or VQ-VAE-based methods (yang2022perceptual; qi2025generative). Figure 5 shows the visual results for the sequence ‘BasketballDrill_832x480_50’ from the HEVC Class C dataset. To ensure a fair comparison, reconstructed sequences from different methods are selected at comparable bitrates. It is observed that the reconstructed videos generated by T-GVC demonstrate superior fidelity in the background regions and key foreground objects (e.g., basketball and the players in red). In contrast, VTM exhibits noticeable blocking artifacts while DCVC-FM suffers from significant high-frequency detail loss, leading to degraded texture quality.
It is worth noting that although some texture and structural details in the reconstructed frames might be partially missing or inaccurate to some extent (e.g., the direction of the ball and the appearance of players in green), the overall visual quality remains satisfactory for the ULB scenarios. Under ULB conditions, we primarily focus on perceptual quality within a given bitrate range. While key semantic information (e.g., the player jumping to catch the ball) is preserved, certain visual artifacts—such as the deblurring effects or minor spatial deviations—are acceptable.
@@ 4.3 Ablation study
To further verify the rationality of our design, we conduct ablation studies on two core components of the T-GVC framework: the guidance mechanism and the motion sampler module. Video sequences characterized by large motions, including ‘ParkScene’, ‘BasketballDrill’, ‘PartyScene’, ‘videoSRC02’ and ‘videoSRC05’ from Class B, Class C and MCL-JCV, are selected for evaluation. The coding efficiency is assessed using the Bjøntegaard Delta rate (BD-rate) (bjontegaard2001calculation), where PSNR is replaced by 1 - LPIPS. Negative BD-rate values indicate bitrate saving for equivalent perceptual quality. We apply DynamiCrafter with empty prompts (i.e., only guided by keyframes) as the comparison anchor in our experiments. The reported BD-rate results are calculated by averaging all test sequences.
Ablation study on the guidance mechanism.
In prior generative video coding methods (zhang2024video; wan2025m3), text is commonly adopted as motion guidance. To demonstrate the superiority of our guidance mechanism, we implement three variants of generative frameworks based on DynamiCrafter: one guided by text (denoted as Text), one guided by trajectories (denoted as Trajectory) and the other by both text and trajectories (denoted as Text+Trajectory). Trajectories are sampled through our proposed sparse motion sampler, while text descriptions are generated from the source video using Hunyuan-Large (sun2024hunyuan) without limitation on length. The comparison results are presented in Table 2, where ”W/o motion bitrate” refer to excluding the bitrate consumed by motion-related information when calculating BPP, while ”W/ motion bitrate” indicate that this bitrate is included. It is observed that the trajectory-guided approach consistently achieves better coding performance under both evaluation conditions. In particular, the perceptual quality improvement brought by trajectory guidance makes up for the additional bitrate overhead, offering a distinct advantage over text-based guidance. When overhead is not considered, the trajectory-guided method still outperforms the text-guided method. Combining text and trajectory guidance may enhance texture on simple-motion sequences. Some visual examples in Figure 6 further demonstrate that trajectory guidance can perform more precise motion control, especially in dynamic areas. Text guidance can improve perceptual quality to some extent and combining text and trajectory guidance may enhance texture on simple-motion sequences. However, in more complex cases, text prompts may introduce artifacts that conflict with trajectory-aligned motion, leading to degraded results. For instance, as shown in the first column of Figure 6, the basketball appears behind the net, which violates the physical plausibility.
@@
Ablation study on the motion sampler module.
To validate the effectiveness of the semantic-aware motion sampling pipeline proposed in T-GVC, we further conduct an ablation study on the motion sampler module. The ablation results are presented in Table 3. Specifically, “Sparse” refers to motion trajectories extracted with our semantic-aware motion sampling pipeline. “Dense” denotes applying dense trajectories as guidance without sampling. “Random” corresponds to randomly selecting sparse trajectories with $k$ keypoints ($K<K_{max}$) for each motion instance. We set the maximum keypoint quantity $K_{max}$ as 15 for the optimal R-D performance. Although dense trajectories can provide more precise motion representation for video reconstruction, such excessive bitrate overhead is intolerable for the video coding task. In contrast, our motion sampling pipeline achieves significant bitrate reduction in motion modeling, while preserving perceptual quality comparable to that of dense trajectory guidance. Compared with a random sampling strategy, our approach thoroughly considers the impact of each motion instance on the video semantics, effectively reducing the bitrate while retaining key motion information.
5 Conclusion
In this paper, we propose T-GVC, a novel Trajectory-Guided Generative Video Coding framework designed for ULB scenarios. Unlike existing methods that rely on domain-specific priors or high-level text guidance, T-GVC introduces a semantic-aware sparse motion sampling strategy, extracting trajectories based on semantic importance to retain key temporal information with minimal bitrate overhead. To further enhance motion realism, a training-free guidance mechanism is integrated into the diffusion process, enabling physically plausible reconstructions. Experimental results demonstrate that T-GVC significantly outperforms traditional codecs and state-of-the-art generative methods, especially in terms of motion control precision.
6 Acknowledgments
This work was supported in part by the National Key R&D Program of China (2023YFA1008500), the National Natural Science Foundation of China (NSFC) under grants U22B2035 and 62502116, and China Post-Doctoral Science Foundation under Grant 2025M774315.
Appendix A Theory
Classifier guidance.
Classifier guidance (dhariwal2021diffusion) is a widely-used technique that modifies the sampling process of a pre-trained diffusion model. The primary objective of conditional generative models is to synthesize a target content $z$ given a conditioning signal $c$, ensuring that $z$ aligns with $c$. To formalize the conditional generation process, the target probability distribution $p\left(\left.z\right|c\right)$ can be derived through Bayes formula:
$p\left(\left.z\right|c\right)=\frac{p\left(\left.c\right|z\right)p\left(z\right)}{p\left(c\right)}$ (7)
where $p\left(z\right)$ is unconditional probability distribution and $p\left(\left.c\right|z\right)$ is the probability distribution of the condition $c$ given a certain content $z$, while the prior probability $p\left(c\right)$ is independent of $z$.
And taking the gradient of the log likelihood of both sides of the Equation 7:
$\mathrm{\nabla}_{z}log{\left(p\left(\left.z\right|c\right)\right)}=\mathrm{\nabla}_{z}log{\left(p\left(z\right)\right)}+\mathrm{\nabla}_{z}log{\left(p\left(\left.c\right|z\right)\right)}$ (8)
Note that the denoising process in diffusion has a relationship with the gradient which represents score in Stochastic Differential Equation (SDE) (song2020score):
$\epsilon_{\theta}\left(z_{t},t\right)=-\sqrt{1-\alpha_{t}}\mathrm{\nabla}_{z_{t}}\log{\left(p\left(z_{t}\right)\right)}$ (9)
$\epsilon_{\theta}\left(z_{t},t,c\right)=-\sqrt{1-\alpha_{t}}\mathrm{\nabla}_{z_{t}}\log{\left(p\left(\left.z_{t}\right|c\right)\right)}$ (10)
We could further rewrite Equation 8 as classification-guided sampling:
$\epsilon_{\theta}\left(z_{t},t,c\right)=\epsilon_{\theta}\left(z_{t},t\right)-\sqrt{1-\alpha_{t}}\mathrm{\nabla}_{z_{t}}\log{\left(p\left({\left.c\right|z}_{t}\right)\right)}$ (11)
where $\sqrt{1-\alpha_{t}}\mathrm{\nabla}_{z_{t}}\log{\left(p\left({\left.c\right|z}_{t}\right)\right)}$ guides the sampling process.
Bansal et al. (bansal2023universal) extended Equation 11 to universal guidance as
${\hat{\epsilon}}_{\theta}\left(z_{t},t\right)=\epsilon_{\theta}\left(z_{t},t\right)+s\left(t\right)\cdot{\nabla_{z_{t}}\ell\left(c,f(z_{t})\right)}$ (12)
where $s\left(t\right)$ controls the guidance strength and the off-the-shelf guidance function $f\left(\cdot\right)$ and loss function $\ell\left(\cdot\right)$ could be calculated based on the predicted clean latent $z_{0}$. This theoretical foundation makes training-free motion guidance possible and inspires us to further explore a new paradigm in trajectory-based guidance.
Appendix B Technical detail
Motion instantiation.
To differentiate between distinct motion patterns, the raw trajectories are clustered using HDBSCAN (mcinnes2017hdbscan), a hierarchical density-based algorithm particularly effective in handling variable cluster densities and suppressing noise, which is critical for processing real-world videos with irregular motions. In this stage, $T_{i}$ is first transformed into 2D trajectories by excluding points where $V_{t}^{i\ }=0$. Then, we formulate each $T_{i}$ as a spatio-temporal feature vector $f^{i}_{traj}$ as follows:
[…]
By clustering these features, the corresponding trajectories are categorized into different motion instances $\mathcal{T}^{0},\mathcal{T}^{1},...,\mathcal{T}^{n-1}$ and form the corresponding motion masks.
@@ Appendix B Technical detail
Sparse $K_{max}$ = 5 W/o trajectory bitrate -5.94
W/ trajectory bitrate 4.12
$K_{max}$ = 10 W/o trajectory bitrate -13.84
W/ trajectory bitrate -1.86
$\mathbf{\boldsymbol{K_{max}}}$ = 15 W/o trajectory bitrate -16.84
W/ trajectory bitrate -3.06
$K_{max}$ = 20 W/o trajectory bitrate -10.58
W/ trajectory bitrate 5.82
Table 4: Ablation study on the keypoint quantity.
Video generation with variable length.
Figure 7 demonstrates the operational scenario of our variable-length generation scheme in processing a video clip beyond 16 frames, with a theoretical maximum capacity of 30 frames. Specifically, for each video clip with length $L>16$, we perform a dual-stage generation strategy where the trajectories are divided into two segments, with each segment guiding one generation stage separately. Notably, increasing the number of processing stages facilitates extended video generation with enhanced temporal coherence, while linearly escalating computational demands.
Appendix C Ablation study
Ablation study on the keypoint quantity.
Ablation study on the video reconstruction.
In our study, DynamiCrafter (xing2024dynamicrafter) was selected as the base model due to its superior reconstruction quality compared to other interpolation methods assessed in (zhang2024video). We conduct additional experiments by integrating diffusion-based SEINE (chen2023seine) and flow-based Ltx-Video (hacohen2025ltx) into T-GVC framework as base video models.
Comparison results at two different bitrate points are provided in Table 5. As observed, DynamiCrafter outperforms SEINE and Ltx-Video in terms of LPIPS. Notably, since Ltx-Video applies higher-magnitude compression in the spatiotemporal dimensions, it introduces significant positional deviation and reduce the precision of trajectory-to-feature alignment, leading to notable performance degradation. When using compressed keyframes and text as guidance, we observe that Ltx-Video would generate artifacts and unrealistic motion in the reconstruction, which is a common issue shared by other existing generative video models.
Models LPIPS1 $\downarrow$ LPIPS2 $\downarrow$
SEINE 0.5201 0.5467
Ltx-Video 0.4959 0.5201
DynamiCrafter 0.3631 0.4123
Table 5: Ablation study on the video reconstruction.
Models Interval Settings BD-rate (%)
Fixed Interval 16 W/o motion bitrate -30.81
W/ motion bitrate 21.15
Fixed Interval 24 W/o motion bitrate 28.06
W/ motion bitrate 43.69
Ours Adaptive W/o motion bitrate -16.84
@@ Appendix C Ablation study
Table 6: Ablation study on the keyframe selection.
Ablation study on the keyframe selection.
Although this adaptive approach is not the focus of our work and has been validated in prior work (zhang2024video), we include additional ablation study using unguided adaptive keyframe selection as a baseline. As observed, a larger frame interval greatly degrade quality, while smaller one leads to higher reconstruction quality but results in increased overhead. In contrast, our adaptive method effectively balances these two aspects.
Appendix D Limitation
Despite impressive performance under ULB conditions, T-GVC still presents several limitations. In scenes involving occlusion or deformable motion, T-GVC captures motion modes via key point variation. However, in cases involving transformations, such as rotation, texture artifacts may emerge despite the preservation of 2D motion trajectories. For human-centric videos, although T-GVC generally preserves semantic fidelity, occasional visual artifacts such as blurriness or unnatural limb motion might occur due to the higher motion complexity (e.g., the reconstructed sequence “videoSRC23_1920x1080_24” in the supplementary materials). Besides, the high computational demand of dense trajectory tracking and video reconstruction makes real-time application of T-GVC still a challenging task, which is a common challenge for generative codecs.
Figure 8: Comparison of the R-D performance between GLC and T-GVC on the UVG dataset.
Appendix E Discussion
We further compare our method with the recent VQ-VAE-based generative approach GLC-video (qi2025generative) on the UVG dataset (resized to 720p for fair comparison), using DISTS (ding2020dists) as the evaluation metric. The results are shown in the Figure 8, where the performance of GLC-video is reported as in the original paper. Benefiting from a more compact representation of inter-frame dependencies, T-GVC achieves lower bitrate points compared to GLC-video. However, although the highest bitrate point of T-GVC ($\approx$ 0.004 bpp) delivers comparable perceptual quality to the lowest point of GLC-video-VAE ($\approx$0.012 bpp), there remains a noticeable gap between T-GVC and GLC-video. Since GLC-video employs a VQ-VAE-based image encoder that has been shown to outperform MS-ILLM (muckley2023improving) used in T-GVC under ULB conditions, there is potential for T-GVC to achieve superior reconstruction quality at even lower bitrates through further optimization of its image encoder and the design of more refined guidance strategies.